You are given that the equilibrium price and quantity of a product A are $30 and 2kg respectively. What is the marginal revenue when the price elasticity of demand is -3?
What is the marginal revenue when quantity is 4?
1
Expert's answer
2020-04-07T08:21:58-0400

Solution:


"MR=p\times (1+ \frac {1}{E_d})"

"MR=30(1-\frac {1}{3})"


"MR=20"

"E_d=Q_d^{/}\times \frac {p}{Q}"

"-3= Q_d^{/}\times \frac {30}{2}"

"Q_d^{\/}=-0.2"

If q=4


"-3=-0.2 \times \frac{p}{4}"

"p=60"

"MR=60(1-\frac {1}{3})"

"MR=40"